Let y be the solution to the following initial value problem.

dy/dt= 4/1+t^2 -2te^2^t , y(1)= pi-e+1
Answer #1

Here Initial valve problem is dy 26 4 **?- 22 Y CIJ = t-et1 9 Now, 2 at de = 4 + + 2. at 9 dy = (4 + t² - ate at Integeabsing